OpenStudy (anonymous):

-8 divided by (-1/2)

OpenStudy (anonymous):

@jdoe0001 @Whitemonsterbunny17

OpenStudy (anonymous):

@brittanydosey @NormaValenzuela

OpenStudy (anonymous):

@TheSmartOne @EclipsedStar @geny55 @CallMeKiki @hhelpplzzzz

OpenStudy (anonymous):

@surjithayer @gabylovesu @urbanmorgans @yeye77 @Truffles @vivalakoda @KittySoftPaws

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Put it into your calculator and if you do not have a calculator use Google :)

TheSmartOne (thesmartone):

Hint: \(\sf\Huge\frac{a}{\frac{b}{c}}=a \div \frac{b}{c}=a\times \frac{c}{b}=\frac{ac}{b}\)

OpenStudy (anonymous):

but i do not know how to divide negative fractions... :/

jimthompson5910 (jim_thompson5910):

\[\Large -8 \div -\frac{1}{2}\] \[\Large -\frac{8}{1} \div -\frac{1}{2}\] \[\Large -\frac{8}{1} \times -\frac{2}{1} = ??\] notice how -8 turned into -8/1 also how I flipped the second fraction and multiplied

OpenStudy (anonymous):

Well the answer comes out to be 16 if that is what you are looking for :) I am too tired to explain it right now :P

OpenStudy (anonymous):

16

OpenStudy (anonymous):

i thought it was - 16?

OpenStudy (anonymous):

no it's 16 no "-"

OpenStudy (anonymous):

But now i know it is 16 becuase i just used my calculator

OpenStudy (anonymous):

THX EVERYONE YOU ALL DESERVE A MEDAL! :3

OpenStudy (anonymous):

i know that's what i did

TheSmartOne (thesmartone):

A negative number times a negative number is postive.

OpenStudy (anonymous):

so can everyone give everyone a medal? :3

TheSmartOne (thesmartone):

positive*

OpenStudy (anonymous):

Bye everyone i am going to close this question!

OpenStudy (samsungfanboy):

:D

OpenStudy (anonymous):

bye XD see ya later